OpenStudy (anonymous):

Medal! I need help with B and C. There are 20 parrots at the animal sanctuary. Their population is increasing at a rate of 15% per year. There are also 24 snakes at the sanctuary. Each year 4 more snakes are born. Part B: How many parrots are at the sanctuary after 10 years? How many snakes are at the sanctuary after the same number of years? Assume there are no deaths to the animals during this time. Part C: After approximately how many years is the number of parrots and snakes the same? Justify your answer mathematically.

OpenStudy (alivejeremy):

Part A) The parrot population: y=20(1.15)x The snake population: y=24+4x Part B) How many parrots are at the sanctuary after 10 years? y=20(1.15)10=80.9≈81 parrots How many snakes are at the sanctuary after the same number of years? 24+4(10)=64 snakes Part C: After approximately how many years is the number of parrots and snakes the same? Justify your answer mathematically. When the two curves intersect , y = y and x = x. We can substitute for the y's and solve the equation: 20(1.15)x=24+4x x≈=6.63025 years

OpenStudy (anonymous):

i seen that one

OpenStudy (anonymous):

Part B: In order to solve for the number of parrots, simply multiply \[20 \times 15%\]. That will give you the number of parrots born each year. Multiply that number by 10 and add it to the original 20 in order to get the total number of parrots. In order to calculate the number of snakes, multiply the increase (4) by the years (10) and add that to the original number of snakes (24)

OpenStudy (alivejeremy):

its right tho

OpenStudy (alivejeremy):

but if it don't help ok

OpenStudy (anonymous):

my teacher would know that